OpenStudy (anonymous):

Find probability of drawing a face card and then a ten from a standard 52 card deck of cards without replacement.

OpenStudy (anonymous):

There are jacks, queens, kings. So you have 3 face cards. 4 suits, thats 12 face cards. So you have 12/52 for a face card. Then without replacing it you have 51 cards left. There are 4 10's so you have 4/51. So you have (12/52)(4/51)=48/2652=12/663

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem :P

OpenStudy (anonymous):

10 isn't a face card? lol

myininaya (myininaya):

is ace a face card?

OpenStudy (anonymous):

I'm not sure, I've never treated it as such.

OpenStudy (anonymous):

isnt an ace a face nvm myininaya beat me to it. lol

OpenStudy (anonymous):

its not a number card for sure, its gotta be counted as a face card.

myininaya (myininaya):

lol im not sure what the ace card is considered

OpenStudy (anonymous):

Isn't it considered as "one"

myininaya (myininaya):

i guess not since it doesn't have a face on it like king , jack , queen

OpenStudy (anonymous):

wikipedia says its not lol

OpenStudy (anonymous):

Jeez, just jump down my throat. I KNOW MY CARDS! lol